### v1

# Jacobian-Enhanced Neural Networks

Abstract
Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. Their main benefit is better accuracy with fewer training points compared to standard neural networks. These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models. Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization. This work derives the complete theory and exemplifies its superiority over standard neural nets for surrogate-based optimization.
Index Terms:
Jacobian, neural network, gradient-enhanced, multi-layer perceptron, surrogate-based optimization, surrogate model, meta-model.Nomenclature
| number of layers in neural network | |
| activation output, | |
| bias (parameter to be learned) | |
| activation function | |
| number of training examples | |
| number of nodes in a layer | |
| number of nodes in input layer, | |
| number of nodes in output layer, | |
| weight (parameter to be learned) | |
| training data input | |
| training data output | |
| predicted output | |
| linear activation input | |
| Hyperparameters | |
| learning rate | |
| hyperparameter to priorities training data values | |
| hyperparameter to priorities training data partials | |
| hyperparameter used for regularization | |
| Mathematical Symbols and Operators | |
| cost function | |
| loss function | |
| first derivative, | |
| second derivative, | |
| Superscripts/Subscripts | |
| node in layer, | |
| input, , | |
| output, , | |
| layer of neural network, | |
| node of current layer, | |
| node of previous layer, | |
| training example, |
I Introduction
Jacobian-Enhanced Neural Networks (JENN) are fully connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of a gradient for vector functions with more than one output. The main benefit of jacobian-enhancement (synonymous to gradient-enhancement) is better accuracy with fewer training points compared to standard fully connected neural nets. This paper develops the theory behind JENN and quantifies its benefits for the intended use-case. It complements previous work in the following ways. First, it provides a concise but complete derivation, including equations in vectorized form (see appendix), without relying on automatic differentiation. Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs. Finally, through judicious use of hyperparameters, it enables weighting; i.e. a process by which individual sample points can be prioritized differently in situations that call for it. The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve SBO in gradient-descent applications.
I-A Intended Use-Case
JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models. Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. This opens the door to useful design applications, such as Monte Carlo simulations of computationally expensive functions for example. However, this is true for any type of surrogate model in general. In the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: Surrogate-Based Optimization (SBO). The field of aerospace engineering is rich in design applications of such a use-case, with entire book chapters dedicated to it [1, 2] and applied examples commonly found in aerodynamic applications [3, 4, 5].
I-B Limitations
First and foremost, gradient-enhanced methods require factors and responses to be smooth and continuous. Second, the value proposition of a surrogate is that the computational expense of generating training data to fit the model is much less than the computational expense of directly performing the analysis with the original model itself, which is not always true, especially when computing partials. Gradient-enhanced methods are therefore only beneficial if the cost of obtaining partials is not excessive in the first place (e.g. adjoint methods [6] in computational fluid dynamics), or if the need for accuracy outweighs the cost of computing the partials. One should therefore carefully weigh the benefit of gradient-enhanced methods relative to the needs of their application.
I-C Supporting Software
An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/. There exist many excellent deep learning frameworks, such as tensorflow [7], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [8]. However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, and requires non-trivial effort to be implemented in those tools. The present library seeks to close that gap; it is intended for those engineers in a rush with a need to predict partials accurately and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [9] is JENN with a different API (present author contribution).
II Previous Work
A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes. Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient Python libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [9], and even review papers about gradient-enhanced meta-models do not mention neural networks, except in passing [10]. However, early work in the late 1990’s can be found [11] which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization. Those authors conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
Follow-on research by that group provides the mathematical derivation of the equations for neural net gradient-enhancement [12], but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility. This shortcoming is overcome in the late 2000’s by a deep, neural-network formulation with gradient-enhancement [3], in which the authors apply their algorithm to surrogate-based optimization of airfoil designs using evolutionary algorithms. All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
Recent work in 2019 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools [4]. Those authors successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not counter-productive in the first place. While deep learning frameworks have obvious advantages (e.g. parallel computing), there is complementary research value in developing the equations for back-propagation, which is a special case of automatic differentiation for scalar functions. Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [13] (PINN). This creative research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator). However, PINN is outside the target scope of the present work, which seeks to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
III Mathematical Derivation
Jacobian-Enhanced Neural Networks (JENN) follow the same steps as standard Neural Networks (NN), shown in Fig. 1, except that the cost function used during parameter update is modified to account for partial derivatives, with the consequence that back propagation must be modified accordingly. This section provides the mathematical derivation for Jacobian-enhancement.
III-A Notation
For clarity, let us adopt a similar notation to the one used by Andrew Ng [14] where superscript denotes the layer number, is the training example, and subscript indicates the node number in a layer. For instance, the activation of the node in layer evaluated at example would be denoted:
To avoid clutter in the derivation to follow, it will be understood that subscripts and refer to quantities associated with the current and previous layer, respectively, evaluated at example without the need for superscripts. Concretely:
Derivatives with respect to input are denoted by the prime symbol and subscript . For example:
In turn, vector quantities will be defined in bold font. For instance, the input and output layers associated with example would be given by:
III-B Normalize Training Data
Normalizing the training data can greatly improve performance when either inputs or outputs differ by orders of magnitude. A simple way to improve numerical conditioning is to subtract the mean and divide by the standard deviation of the data. The normalized data then becomes:
This approach works well, except when all values are very close such that the denominator vanishes as standard deviation goes to zero. In that case, alternative normalization methods (or none) should be considered.
III-C Forward Propagation
Predictions are obtained by propagating information forward throughout the neural network in a recursive manner, starting with the input layer and ending with the output layer. For each layer (excluding input layer), activations are computed as:
where are neural network parameters to be learned. For the hidden layers (), the activation function is taken to be the hyperbolic tangent222Other choices are possible, provided the function is smooth (i.e. not ReLu):
For the output layer (), the activation function is linear:
III-D Parameter Update
The neural network parameters are learned by solving the following, unconstrained optimization problem:
where it should be understood that is short hand for all parameters of the neural network. The cost function is given by the following expression, where is a hyper-parameter to be tuned which controls regularization:
Parameters are updated iteratively using gradient-descent:
where the hyper-parameter is the learning rate. In practice, some improved version of gradient-descent would be used instead, such ADAM [15], but the former is easier to explain without loss of generality. The loss function is taken to be a modified version of the Least Squares Estimator (LSE), augmented with an extra term to account for partial derivatives:
Note that two more hyper-parameters have been introduced: and . They act as weights that can be used to prioritize subsets of the training data in situations that call for it. By default, , meaning that all data is prioritized equally, which is the common use-case. However, in some situations, there may be an incentive to prioritize certain examples over others, as will be shown in the results section. This formulation offers the flexibility to do so.
III-E Backward Propagation
Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.r.t.) the inputs . By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
III-E1 Loss Function Derivatives w.r.t. Neural Net Parameters
In a first step, assume the quantities and are known and consider the gradient of the loss function with respect to the parameters from layer . Applying the chain rule to Eq. III-D yields:
where the superscript has been dropped for notational convenience, with the understanding that all equations to follow apply to example . The various partial derivatives in Eq. 11 follow from Eq. 3:
Substituting these expressions into Eq. 11 yields:
III-E2 Loss Function Derivatives w.r.t. to Neural Net Activations
All quantities in Eq. 15 are known, except for and which must be obtained recursively by propagating information backwards from one layer to the next, starting with the output layer . From Eq. III-D, it follows that:
The derivatives for the next layer are obtained through the chain rule. According to Eq. III-D, the loss function depends on activations and associated partials , which are themselves functions of and . Hence, according to the chain rule, current layer derivatives w.r.t. to previous layer activations are:
Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds that:
All necessary equations to compute the derivatives of the loss function w.r.t. parameters in any layer are now known. Derivatives are obtained by working our way back recursively across layers, repeatedly applying Eq. 17 after passing the solution in the current layer to the preceding layer as follows:
III-E3 Cost Function Derivatives w.r.t. Parameters
IV Verification and Validation
The theory developed in the previous section was implemented into an installable Python library for which a link was provided earlier. This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig. 2, which compares JENN against a standard Neural Net (NN) when training data is sparse (shown as dots). Higher dimensional examples are available on the project’s homepage, but were not included since it is hard to visualize more than two dimensions on a static plot. All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data (provided the cost of obtaining the Jacobian is not counter-productive). Finally, it was also verified that runtime scales as using the Rastrigin function to generate progressively larger datasets and training a new model, as shown in Fig. 2c using a last generation MacBook Pro (2021).
V Example Application: Surrogate-Based Optimization
Surrogate-based optimization is the process of using a fast-running approximation, such as a neural net, to optimize some slow-running function of interest, rather than directly using the function itself. This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application. For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory[16]. The purpose of this section is to quantify the benefit of JENN relative to regular neural networks for gradient-based optimization. This will be accomplished by solving the following optimization problem:
| Minimize: | (21) | |||
| With Respect To: | (22) | |||
| Subject To: | (23) | |||
where is a surrogate-model approximation to the famous Rosenbrock function:
Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3 where optimizer convergence histories are shown for the true response and its surrogate-model approximations using JENN and NN under the same settings, except for jacobian-enhancement, respectively. Training data (not shown to avoid clutter) was generate by sampling the true response on a regular grid. Training histories are provided in Figs. 3a–3b to verify that both neural networks converged to the same degree.
V-A Results Without Polishing
In a first step, consider Fig.3c which does not use polishing (explained in the next section). Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed. For both surrogates, the terminal value is indeed close to zero (the true optimum) but neither JENN nor NN reach the true location of this optimum situated at ; although JENN does get much closer. The Rosenbrock function is notorious for exhibiting a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization. Carefully studying the contour lines, one can indeed observe subtle undulations in this shallow valley within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
This finding is not surprising. Recall from Eq. III-D that neural nets learn by minimizing prediction error. Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting small features in the vicinity of the optimum, where prediction error implies negigible penalty. This situation is at odds with the needs of gradient-based optimization, which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
V-B Results With Polishing
For most practical engineering problems, close is good enough. However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design (to ensure smooth contours free of undulations). This situation can be ameliorated by using what will be called polishing. It consists of restarting a previously trained model and continue training it after magnifying some portion of the design space to be improved. This is accomplished by judiciously setting or to prioritize specific training points belonging to that subspace. For example, a radial basis function centered on zero could be used to magnify small numbers in the aforementioned shallow valley of the Rosenbrock function:
where and are hyperparameters that control the radius of influence of a point. They were set to and in the results shown. Far away from any point where , the weights are , whereas nears those points . This has the desired effect of disproportionately emphasizing small values of and during training by up to a factor of 1000. The results are shown in Fig. 3d where it can be seen that JENN now recovers the true optimum almost exactly; artificial local minima nearly vanish. This is not true for NN; although results improve noticeably. Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
VI Noisy Partials
We now consider the effect of noisy partials. As mentioned in the introduction, many practical design applications rely on legacy software suites, such as missile DATCOM [17] during early design phases, for which analytical or automatically differentiated partials are not available. In this case, the only recourse for computing derivatives is finite difference. However, despite their attractive simplicity, it has been shown that finite difference methods are not as accurate as their counterparts [18]. It is therefore natural to wonder: how accurate must partials be in order to take advantage of JENN in design applications where finite difference is unavoidable?
The answer to this question is shown by example in Fig. 4 using the Rastrigin function as a test bed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural net (gradient-enhancement off). Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes. The corresponding mean finite difference error shown was computed against true analytical partials. It can clearly be seen that beyond seven percent error, the benefit of gradient-enhancement vanishes. In fact, beyond that threshold, partials are too inaccurate to provide useful information: accounting for erroneous partials is worse than not accounting for partials at all. These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
Finally, these results further show that there is no benefit to setting to values other than zero or one; it does not mitigate the effect of noisy partials. For most applications, the hyperparameter should therefore be treated as a binary scalar to turn gradient-enhancement on () or off (); except in special situations which call for distribution (rather than a scalar) such as polishing, as previously illustrated, or incomplete partials. The latter is likely to be a common scenario where partials are not available for all inputs, in which case can be used to turn gradient-enhancement on/off depending on whether a partial is missing or not (i.e. resulting in a tensor of zeros and ones).
VII Conclusion
In summary, this research developed a jacobian-enhanced, deep-learning formulation for dense multi-layer perceptrons and successfully validated the implementation against selected canonical test functions. It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place. It was further shown that JENN outperforms standard neural nets in the context of surrogate-based optimization, getting much closer to the true optimum but not quite reaching it for the selected example. In fact, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero. However, it was shown that this shortcoming can be overcome through a process termed polishing, which prioritizes salient sample points in that neighborhood through judicious use of hyperparameters. This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it. The open-source Python library associated with this work was provided as a link in earlier sections, in the hope that it will be useful to others and spark new contributions in this researcher area.
Acknowledgement
The supporting code for this work used the exercises by Prof. Andrew Ng [14] as a starting point. It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied). The author would therefore like to thank Andrew Ng for offering the fundamentals of deep learning on Coursera, which took a complicated subject and explained it in simple terms that even an aerospace engineer could understand.
VIII Key Equations in Vectorized Form
Key equations from the preceding material will now be re-written in a vectorized form (i.e. processing all examples simultaneously without loops) suitable for programming, as a helpful companion to source code. In the notation to follow, the symbol implies element-wise array multiplication.
VIII-A Data Structures
Using capital letters, notation can be extended to matrix notation to account for all training examples, starting with training data inputs and outputs:
The associated Jacobians are formatted as:
Note that the hyperparameters can either be scalars or real arrays of the same shape as and , respectively. The shape of the neural network parameters to be learned (i.e. weights and biases) is independent of the number of training examples. They depend only on layer sizes and are denoted as:
However, hidden layer activations do depend on the number of examples and are stored as:
The associated hidden layer derivatives w.r.t. inputs are stored as:
It follows that hidden layer derivatives w.r.t. a specific input are formatted as:
Finally, the back propagation derivatives w.r.t. parameters are:
VIII-B Forward Propagation (Eq. 3)
VIII-C Parameter Update (Eq. 8)
VIII-D Backward Propagation (Eqs. 15–16, 17, 20)
The loss function partials w.r.t. last layer activations are:
The loss function partials w.r.t. hidden layer activations are:
The loss function partials w.r.t. neural net parameters are:
Finally, the cost function partials w.r.t. neural net parameters are:
where S is used to denote the vectorized summation along the last axis of an array666https://numpy.org/doc/stable/reference/generated/numpy.sum.html. Any other summation shown above implies that it could not be vectorized and required a loop. This only happened when summing over partials. Hence, the number of loops needed equals the number of inputs , which is usually negligible compared to the size of the training data . The latter is the critical dimension to be vectorized.

### v2

@@
 These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models.
-Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly.
+Since a surrogate emulates the original model accurately in real time, it yields a computational speed benefit that can be used to carry out orders of magnitude more function calls quickly.
 However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization.
@@
 Superscripts/Subscripts | |
-| node in layer, | |
+node in current layer, | |
 input, , | |
@@
 layer of neural network, | |
+node of previous layer, | |
 node of current layer, | |
-| node of previous layer, | |
 training example, |
@@
 a process by which individual sample points can be prioritized differently in situations that call for it.
-The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve SBO in gradient-descent applications.
+The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve Surrogate-Based Optimization (SBO) in gradient-descent applications.
 I-A Intended Use-Case
 JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models.
-Since a surrogate emulates the original model accurately in near-real time, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly.
-This opens the door to useful design applications, such as Monte Carlo simulations of computationally expensive functions for example.
-However, this is true for any type of surrogate model in general.
-In the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: Surrogate-Based Optimization (SBO).
-The field of aerospace engineering is rich in design applications of such a use-case, with entire book chapters dedicated to it [1, 2] and applied examples commonly found in aerodynamic applications [3, 4, 5].
+Since a surrogate emulates the original model accurately in fractions of second, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly.
+This allows designers to explore a broader tradespace, and more readily quantify the impact of various parameters in order to better understand the performance of the system under design.
+While this is true of any surrogate modeling approach, JENN has an additional benefit that most surrogate modeling approaches lack: the ability to predict partial derivatives with improved accuracy.
+This is a critical property for one important use-case: surrogate-based optimization.
+The field of engineering is rich in design applications of such a use-case with entire book chapters dedicated to it [1, 2, 3, 4], review papers [5, 6, 7, 8], and countless examples to be found, particularly in the field of aerospace engineering [9, 10, 11, 12].
 I-B Limitations
@@
 An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/.
-There exist many excellent deep learning frameworks, such as tensorflow [7], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [8].
-However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, and requires non-trivial effort to be implemented in those tools.
-The present library seeks to close that gap; it is intended for those engineers in a rush with a need to predict partials accurately and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [9] is JENN with a different API (present author contribution).
+There exist many excellent deep learning frameworks, such as TensorFlow [14] and PyTorch [15], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [16].
+However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, which requires non-trivial effort to be implemented in those tools.
+The present library seeks to close that gap.
+It is intended for engineers with limited time, who have a need to predict partials accurately, and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [17] is JENN with a different API (present author contribution).
 II Previous Work
 A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes.
-Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient Python libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [9], and even review papers about gradient-enhanced meta-models do not mention neural networks, except in passing [10].
-However, early work in the late 1990’s can be found [11] which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization.
-Those authors conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
-Follow-on research by that group provides the mathematical derivation of the equations for neural net gradient-enhancement [12], but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility.
-This shortcoming is overcome in the late 2000’s by a deep, neural-network formulation with gradient-enhancement [3], in which the authors apply their algorithm to surrogate-based optimization of airfoil designs using evolutionary algorithms.
+Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient open-source libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [17], and even review papers about gradient-enhanced meta-models only mention neural networks in passing [18].
+However, early work by Sellar and Batill [19] can be found in the late 1990’s, which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization.
+They conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
+Follow-on research by that Liu and Batill [20] provides the mathematical derivation of the equations for neural net gradient-enhancement, but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility.
+This shortcoming is overcome by Giannakoglou et al.
+[10] in the late 2000’s who developed a deep neural-network formulation with gradient-enhancement and use it for aerodynamic design applications.
 All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
-Recent work in 2019 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools [4].
-Those authors successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not counter-productive in the first place.
+Recent work by Bouhlel et al.
+[11] in 2020 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools.
+Bouhlel et al.
+successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not prohibitive in the first place.
 While deep learning frameworks have obvious advantages (e.g.
@@
 Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
+The works surveyed so far have focused on prediction accuracy, but Hoffman et al.
+[21] take a different approach by focusing on robustness instead.
+They introduce the concept of Jacobian regularization for classification, where the key idea is to ensure robustness of models against input perturbations due to adversarial attacks; for example, a purposefully modified input image to avoid detection.
+However, robustness is outside the scope of the present work, which is to predict design partials accurately given high quality inputs.
 Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [22] (PINN).
-This creative research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator).
-However, PINN is outside the target scope of the present work, which seeks to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
+This novel research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator).
+However, PINN is outside the scope of the present work, which is to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
 III Mathematical Derivation
@@
 where the hyper-parameter is the learning rate.
-In practice, some improved version of gradient-descent would be used instead, such ADAM [15], but the former is easier to explain without loss of generality.
-The loss function is taken to be a modified version of the Least Squares Estimator (LSE), augmented with an extra term to account for partial derivatives:
+In practice, some improved version of gradient-descent would be used instead, such as ADAM [24], but the former is easier to explain without loss of generality.
+The loss function is taken to be a modified version of the Least rquares Estimator (LSE), augmented with an extra term to account for partial derivatives:
 Note that two more hyper-parameters have been introduced: and .
@@
 III-E Backward Propagation
-Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.r.t.) the inputs .
+Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.s.t.) the inputs .
 By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
-III-E1 Loss Function Derivatives w.r.t. Neural Net Parameters
+III-E1 Loss Function Derivatives w.s.t. Neural Net Parameters
 In a first step, assume the quantities and are known and consider the gradient of the loss function with respect to the parameters from layer . Applying the chain rule to Eq. III-D yields:
 where the superscript has been dropped for notational convenience, with the understanding that all equations to follow apply to example . The various partial derivatives in Eq. 11 follow from Eq. 3:
-Substituting these expressions into Eq. 11 yields:
-III-E2 Loss Function Derivatives w.r.t. to Neural Net Activations
+rubstituting these expressions into Eq. 11 yields:
+III-E2 Loss Function Derivatives w.s.t. to Neural Net Activations
 All quantities in Eq.
@@
 III-D, the loss function depends on activations and associated partials , which are themselves functions of and .
-Hence, according to the chain rule, current layer derivatives w.r.t.
+Hence, according to the chain rule, current layer derivatives w.s.t.
 to previous layer activations are:
 Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds that:
-All necessary equations to compute the derivatives of the loss function w.r.t.
+All necessary equations to compute the derivatives of the loss function w.s.t.
 parameters in any layer are now known.
@@
 17 after passing the solution in the current layer to the preceding layer as follows:
-III-E3 Cost Function Derivatives w.r.t. Parameters
+III-E3 Cost Function Derivatives w.s.t. Parameters
 IV Verification and Validation
@@
 This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig.
-2, which compares JENN against a standard Neural Net (NN) when training data is sparse (shown as dots).
-Higher dimensional examples are available on the project’s homepage, but were not included since it is hard to visualize more than two dimensions on a static plot.
-All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data (provided the cost of obtaining the Jacobian is not counter-productive).
-Finally, it was also verified that runtime scales as using the Rastrigin function to generate progressively larger datasets and training a new model, as shown in Fig.
-2c using a last generation MacBook Pro (2021).
+2, where training data is shown as dots.
+These validation results compare JENN against a standard Neural Net (NN), defined as the equivalent neural network without gradient-enhancement.
+All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data, so long as the computational cost of obtaining the Jacobian is not prohibitive in the first place.
+In Figs.
+2a and 2b, where there are only three and four training samples respectively, NN fails to even capture the correct trend whereas JENN predicts the response almost perfectly.
+In Fig.
+2d, where there are 100 samples, NN correctly captures the trend but contour lines are distorted, implying more samples would be needed to improve prediction.
+By contrast, under the same conditions, JENN predicts the Rastrigin function [25] almost perfectly.
+Finally, it was also verified that runtime scales as , as shown in Fig.
+2c.
+This plot was obtained using the Rastrigin function to generate progressively larger datasets, train a new model, and record runtime on a MacBook Pro (2021).
 V Example Application: Surrogate-Based Optimization
-Surrogate-based optimization is the process of using a fast-running approximation, such as a neural net, to optimize some slow-running function of interest, rather than directly using the function itself.
+Surrogate-based optimization is the process of using a fast-running approximation, such as a neural networks, to optimize some slow-running function of interest, rather than directly using the function itself.
 This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application.
@@
 Subject To: | (23) | |||
-where is a surrogate-model approximation to the famous Rosenbrock function:
-Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3 where optimizer convergence histories are shown for the true response and its surrogate-model approximations using JENN and NN under the same settings, except for jacobian-enhancement, respectively.
-Training data (not shown to avoid clutter) was generate by sampling the true response on a regular grid.
-Training histories are provided in Figs.
-3a–3b to verify that both neural networks converged to the same degree.
+where is a surrogate-model approximation to the Rosenbrock [27] function, which is used as an optimization benchmark because it exhibits a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization:
+Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3b, where optimization histories are shown for the true response and three surrogate-model approximations: NN, JENN, and JENN (polished).
+Training data is not shown to avoid clutter, but was generated by sampling the true response on a regular grid plus 100 latin hypercube points.
+The neural network training histories are also included in Fig.
+3a, in order to verify that both neural networks converged to the same degree.
 V-A Results Without Polishing
-In a first step, consider Fig.3c which does not use polishing (explained in the next section).
-Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed.
-For both surrogates, the terminal value is indeed close to zero (the true optimum) but neither JENN nor NN reach the true location of this optimum situated at ; although JENN does get much closer.
-The Rosenbrock function is notorious for exhibiting a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization.
-Carefully studying the contour lines, one can indeed observe subtle undulations in this shallow valley within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
-This finding is not surprising.
-Recall from Eq.
-III-D that neural nets learn by minimizing prediction error.
-Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting small features in the vicinity of the optimum, where prediction error implies negigible penalty.
-This situation is at odds with the needs of gradient-based optimization, which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
+Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed: neither JENN nor NN reach the true location of this optimum situated at .
+Carefully studying the contour lines of NN in the region where the Rosenbrock function flattens out, one can observe subtle undulations, within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
+These effects are less pronounced for JENN, but the optimizer gets trapped nonetheless.
+In hindsight, this finding is not surprising because neural networks learn by minimizing prediction error (recall Eq.
+III-D).
+Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting flat regions in the vicinity of a zero optimum, where prediction error penalties are negligible.
+This situation is at odds with the needs of gradient-based optimization which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
+For a broad set of engineering problems, the results attained by these baseline models may be sufficiently accurate.
+However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design where it is imperative that the final airfoil is free of undulations and non-smooth contours.
+An imperative research question therefore arises: how can these competing needs be reconciled?
 V-B Results With Polishing
-For most practical engineering problems, close is good enough.
-However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design (to ensure smooth contours free of undulations).
-This situation can be ameliorated by using what will be called polishing.
-It consists of restarting a previously trained model and continue training it after magnifying some portion of the design space to be improved.
-This is accomplished by judiciously setting or to prioritize specific training points belonging to that subspace.
-For example, a radial basis function centered on zero could be used to magnify small numbers in the aforementioned shallow valley of the Rosenbrock function:
+This situation can be ameliorated through “polishing.” Given a trained model, this approach consists of train it further after magnifying regions of interest.
+This is accomplished by judiciously setting to prioritize specific training points belonging to that subspace.
+For example, flat regions could be magnified by allocating more importance to small slopes during training, using a radial basis functions centered on such as:
 where and are hyperparameters that control the radius of influence of a point.
 They were set to and in the results shown.
-Far away from any point where , the weights are , whereas nears those points .
-This has the desired effect of disproportionately emphasizing small values of and during training by up to a factor of 1000.
-The results are shown in Fig.
-3d where it can be seen that JENN now recovers the true optimum almost exactly; artificial local minima nearly vanish.
-This is not true for NN; although results improve noticeably.
+Far away from any point where , the weights are whereas near those points .
+This has the desired effect of disproportionately emphasizing regions of small slope during training by up to a factor of 1000 at the expense of de-emphasizing other regions, which is why it is necessary to start from an already trained model.
+Final results are shown in Fig.
+3b for the Rosenbrock test case, where it can be seen that JENN (polished) now recovers the true optimum almost exactly.
 Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
@@
 The answer to this question is shown by example in Fig.
-4 using the Rastrigin function as a test bed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural net (gradient-enhancement off).
+4 using the Rastrigin function as a testbed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural network (gradient-enhancement off).
 Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes.
@@
 It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place.
-It was further shown that JENN outperforms standard neural nets in the context of surrogate-based optimization, getting much closer to the true optimum but not quite reaching it for the selected example.
-In fact, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero.
-However, it was shown that this shortcoming can be overcome through a process termed polishing, which prioritizes salient sample points in that neighborhood through judicious use of hyperparameters.
+It was further shown that JENN outperforms standard neural networks in the context of surrogate-based optimization.
+However, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero.
+It was further shown that this shortcoming can be overcome through polishing, which can be used to prioritize salient regions of interest through judicious use of hyperparameters.
 This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it.
-The open-source Python library associated with this work was provided as a link in earlier sections, in the hope that it will be useful to others and spark new contributions in this researcher area.
+The ideas and findings of this paper have been published as an open-source library to allow readers to validate the results, and hopefully assist them in their surrogate modeling efforts.
 Acknowledgement
-The supporting code for this work used the exercises by Prof.
-Andrew Ng [14] as a starting point.
+The supporting code for this work used the exercises by Professor Andrew Ng [23] as a starting point.
 It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied).
-The author would therefore like to thank Andrew Ng for offering the fundamentals of deep learning on Coursera, which took a complicated subject and explained it in simple terms that even an aerospace engineer could understand.
+The author would therefore like to thank Professor Andrew Ng for offering the fundamentals of deep learning on Coursera, which masterfully explained what can be considered a challenging and complicated subject in an easy to understand form.
 VIII Key Equations in Vectorized Form
-Key equations from the preceding material will now be re-written in a vectorized form (i.e.
-processing all examples simultaneously without loops) suitable for programming, as a helpful companion to source code.
-In the notation to follow, the symbol implies element-wise array multiplication.
+This section provides vectorized equations to facilitate re-implementation or as a companion guide to the JENN library.
 VIII-A Data Structures

### v3

@@
 Abstract
-Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately.
+Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives more accurately.
 Their main benefit is better accuracy with fewer training points compared to standard neural networks.
-These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models.
-Since a surrogate emulates the original model accurately in real time, it yields a computational speed benefit that can be used to carry out orders of magnitude more function calls quickly.
-However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization.
-This work derives the complete theory and exemplifies its superiority over standard neural nets for surrogate-based optimization.
-Index Terms:
-Jacobian, neural network, gradient-enhanced, multi-layer perceptron, surrogate-based optimization, surrogate model, meta-model.Nomenclature
+These attributes are particularly desirable in the field of design optimization, where there is often the need to replace computationally expensive models with fast running approximations, known as surrogate models (or metamodels), to speed up solutions.
+In the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which yields improved accuracy for gradient-based optimization.
+This work develops the theory of Jacobian-Enhanced Neural Nets and applies it to an airfoil optimization problem in order to illustrate its benefits over the status quo.
+keywords:
+Jacobian , gradient-enhanced , neural networks , surrogate model , aerodynamic shape optimization[ELSYS]organization=Sr Researcher Engineer at Georgia Tech Research Institute, city=Atlanta, state=GA, postcode=30332, country=USA
+A neural network formulation capable of predicting partial derivatives more accurately with significantly fewer training data samples
+Practical illustration through applied aerospace engineering example, taken to be aerodynamic shape optimization using CFD
+Complete mathematical derivation of the underlying theory for arbitrary deep network, including vectorized equations for computer programming, with open-source Python code implementation
+Quantification of advantages & limitations through canonical example, including analysis of noisy partials and performance on benchmark optimization problem
+Convenient framework to handle missing partials, believed to be a common use-case of practical importance.
+Nomenclature
 | number of layers in neural network | |
@@
 | training example, |
-I Introduction
-Jacobian-Enhanced Neural Networks (JENN) are fully connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately.
-This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of a gradient for vector functions with more than one output.
-The main benefit of jacobian-enhancement (synonymous to gradient-enhancement) is better accuracy with fewer training points compared to standard fully connected neural nets.
-This paper develops the theory behind JENN and quantifies its benefits for the intended use-case.
-It complements previous work in the following ways.
-First, it provides a concise but complete derivation, including equations in vectorized form (see appendix), without relying on automatic differentiation.
+1 Introduction
+Aerodynamic shape optimization is concerned with maximizing aerodynamic performance by changing the shape of a body subjected to specific flow conditions.
+The state of the art uses computational fluid dynamics and adjoint methods to achieve this goal (Jameson1988; Kenway2019).
+These methods scale very well, enabling optimization problems with hundreds of decision variables to be solved as efficiently as problems with a handful.
+However, the cost of individual function calls remains an impediment: a single flow solution can take several hours depending on the use case.
+For this reason, there has been significant interest in accelerating numerical computations through machine learning in aerodynamic shape optimization (Li2022), and the present work focuses on one method in particular: Jacobian-Enhanced Neural Networks (JENN).
+1.1 Overview
+JENN are fully connected multi-layer perceptrons, whose training process has been modified to predict partial derivatives more accurately.
+This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of the gradient for functions of more than one output.
+The main benefit of Jacobian enhancement is better accuracy with fewer training points compared to standard fully connected neural nets.
+This paper develops the theory behind JENN and applies it to an airfoil optimization problem in order to illustrate its benefits over status quo neural networks.
+It complements previous work on gradient-enhanced neural nets in three ways.
+First, it provides a concise and complete derivation of the theory, including equations in vectorized form, without relying on automatic differentiation.
 Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs.
-Finally, through judicious use of hyperparameters, it enables weighting; i.e.
-a process by which individual sample points can be prioritized differently in situations that call for it.
-The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve Surrogate-Based Optimization (SBO) in gradient-descent applications.
-I-A Intended Use-Case
-JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models.
-Since a surrogate emulates the original model accurately in fractions of second, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly.
-This allows designers to explore a broader tradespace, and more readily quantify the impact of various parameters in order to better understand the performance of the system under design.
-While this is true of any surrogate modeling approach, JENN has an additional benefit that most surrogate modeling approaches lack: the ability to predict partial derivatives with improved accuracy.
-This is a critical property for one important use-case: surrogate-based optimization.
-The field of engineering is rich in design applications of such a use-case with entire book chapters dedicated to it [1, 2, 3, 4], review papers [5, 6, 7, 8], and countless examples to be found, particularly in the field of aerospace engineering [9, 10, 11, 12].
-I-B Limitations
+Third, through judicious use of hyperparameters, it enables individual sample points to be prioritized differently in situations that call for it, the utility of which is illustrated in the application section.
+1.2 Intended Use-Case
+Even though JENN was developed with aerodynamic shape optimization in mind, it is a general method that applies to any engineering problem for which partial derivatives are defined.
+It is intended for the field of computer aided design, where there is often a need to replace computationally expensive simulation models with fast running approximations, commonly referred to as surrogate models or metamodels.
+Since a surrogate emulates the original model with close approximation in a fraction of a second, it yields a speed benefit that enables orders of magnitude more function calls to be carried out quickly.
+This allows designers to explore a tradespace more fully, in order to better understand the impact of design variables on the performance of the system under design.
+While this is true of any surrogate modeling approach, JENN has an additional benefit that most surrogate modeling approaches lack: the ability to predict partial derivatives more accurately.
+This is a desirable property for one use case in particular: surrogate-based optimization.
+The field of engineering is rich in design applications of such a use case with entire book chapters dedicated to it (Keane2005; Forrester2011; Iuliano2016; Martins_Ning_2021), review papers (Queipo2005; haridy2023review; su14073867; Wang2006), and countless examples to be found, particularly in the field of aerospace engineering (Robinson2008; Giannakoglou2006; Bouhlel2020; Nagawkar2022).
+1.3 Limitations
 First and foremost, gradient-enhanced methods require factors and responses to be smooth and continuous.
 Second, the value proposition of a surrogate is that the computational expense of generating training data to fit the model is much less than the computational expense of directly performing the analysis with the original model itself, which is not always true, especially when computing partials.
-Gradient-enhanced methods are therefore only beneficial if the cost of obtaining partials is not excessive in the first place (e.g.
-adjoint methods [13] in computational fluid dynamics), or if the need for accuracy outweighs the cost of computing the partials.
-One should therefore carefully weigh the benefit of gradient-enhanced methods relative to the needs of their application.
-I-C Supporting Software
-An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/.
-There exist many excellent deep learning frameworks, such as TensorFlow [14] and PyTorch [15], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [16].
-However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, which requires non-trivial effort to be implemented in those tools.
-The present library seeks to close that gap.
-It is intended for engineers with limited time, who have a need to predict partials accurately, and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [17] is JENN with a different API (present author contribution).
-II Previous Work
-A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes.
-Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient open-source libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [17], and even review papers about gradient-enhanced meta-models only mention neural networks in passing [18].
-However, early work by Sellar and Batill [19] can be found in the late 1990’s, which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization.
-They conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
-Follow-on research by that Liu and Batill [20] provides the mathematical derivation of the equations for neural net gradient-enhancement, but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility.
-This shortcoming is overcome by Giannakoglou et al.
-[10] in the late 2000’s who developed a deep neural-network formulation with gradient-enhancement and use it for aerodynamic design applications.
+Gradient-enhanced methods are therefore only beneficial if the cost of obtaining partials is not excessive in the first place or if the need for accuracy outweighs the cost of computing the partials; see Martins et al.
+(Martins2013a) for an excellent review of all existing methods for computing the derivatives of computational models.
+2 Previous Work
+A cursory literature research reveals that the field of gradient-enhanced learning is predominantly focused on Gaussian processes.
+Dedicated book chapters can be found about gradient-enhanced Kriging models (Forrester2011), salient open-source libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods (saves2024smt), and even review papers about gradient-enhanced metamodels only mention neural networks in passing (Laurent2019).
+However, early work by (Sellar1996) can be found which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization.
+The authors of that work conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all surveyed work to follow.
+Follow-on research by (Liu2000) further provides the mathematical derivation of the equations for neural network gradient-enhancement, but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility.
+This shortcoming is overcome by (Giannakoglou2006) who developed a deep neural network formulation with gradient enhancement and used it for aerodynamic design applications.
 All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
-Recent work by Bouhlel et al.
-[11] in 2020 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools.
-Bouhlel et al.
-successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not prohibitive in the first place.
+More recent work by (Bouhlel2020) takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic differentiation tools.
+Using TensorFlow (tensorflow2015), gradient-enhanced neural networks are successfully used to accelerate airfoil shape optimization, culminating into webfoil111https://webfoil.engin.umich.edu/; a useful web App for fast airfoil shaping.
+Overall, those authors reaffirm the previous conclusion that gradient-enhanced techniques improve accuracy and reduce the amount of training data required for good fits.
 While deep learning frameworks have obvious advantages (e.g.
-parallel computing), there is complementary research value in developing the equations for back-propagation, which is a special case of automatic differentiation for scalar functions.
-Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
-The works surveyed so far have focused on prediction accuracy, but Hoffman et al.
-[21] take a different approach by focusing on robustness instead.
-They introduce the concept of Jacobian regularization for classification, where the key idea is to ensure robustness of models against input perturbations due to adversarial attacks; for example, a purposefully modified input image to avoid detection.
-However, robustness is outside the scope of the present work, which is to predict design partials accurately given high quality inputs.
-Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [22] (PINN).
-This novel research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator).
-However, PINN is outside the scope of the present work, which is to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
-III Mathematical Derivation
+parallel computing), there is complementary research value in developing the equations for back-propagation from scratch, which is one of the objectives of this paper.
+Stand-alone equations are reproducible without third-party tools and convey helpful insight into the mathematics behind a specific algorithm.
+Backprop is a special case of automatic differentiation for scalar functions.
+The works surveyed so far have focused on prediction accuracy, but a keyword search for Jacobian and neural networks also points to the concept of Jacobian regularization (hoffman2019robust) for classification.
+The key idea is to ensure robustness of models against input perturbations due to adversarial attacks (e.g.
+a purposefully modified input image to avoid detection) by using Jacobian information to reduce sensitivity.
+However, robustness is outside the scope of the present work, which seeks to predict partials accurately for design applications given high quality inputs.
+Finally, the most recent work on gradient-enhanced neural networks focuses on gradient-enhanced Physics-Informed Neural Networks (PINN) (yu2022), which uses the residuals of partial differential equations as the loss function for training neural nets.
+It is a deep learning framework for predicting systems governed by nonlinear partial differential equations, in which the neural network effectively learns to solve a system of equations rather than merely emulating a dataset.
+However, PINN falls outside the scope of the present work, which seeks to develop a general purpose algorithm to fit any smooth function regardless of what it represents, be it physics-informed or not.
+3 Mathematical Derivation
 Jacobian-Enhanced Neural Networks (JENN) follow the same steps as standard Neural Networks (NN), shown in Fig.
 1, except that the cost function used during parameter update is modified to account for partial derivatives, with the consequence that back propagation must be modified accordingly.
-This section provides the mathematical derivation for Jacobian-enhancement.
-III-A Notation
-For clarity, let us adopt a similar notation to the one used by Andrew Ng [23] where superscript denotes the layer number, is the training example, and subscript indicates the node number in a layer.
+This section provides the mathematical derivation.
+3.1 Notation
+For clarity, let us adopt a similar notation to the one used by (Ng2017) where superscript denotes the layer number, is the training example, and subscript indicates the node number in a layer.
 For instance, the activation of the node in layer evaluated at example would be denoted:
@@
 Concretely:
-Derivatives with respect to input are denoted by the prime symbol and subscript . For example:
+Derivatives w.r.t.input are denoted by the prime symbol and subscript , e.g.
 In turn, vector quantities will be defined in bold font. For instance, the input and output layers associated with example would be given by:
-III-B Normalize Training Data
+3.2 Normalize Training Data
 Normalizing the training data can greatly improve performance when either inputs or outputs differ by orders of magnitude.
-A simple way to improve numerical conditioning is to subtract the mean and divide by the standard deviation of the data.
+A simple way to improve numerical conditioning is to subtract the mean and divide by the standard deviation of the data, provided the denominator is not close to zero.
 The normalized data then becomes:
-This approach works well, except when all values are very close such that the denominator vanishes as standard deviation goes to zero.
-In that case, alternative normalization methods (or none) should be considered.
-III-C Forward Propagation
+3.3 Forward Propagation
 Predictions are obtained by propagating information forward throughout the neural network in a recursive manner, starting with the input layer and ending with the output layer.
@@
 For the output layer (), the activation function is linear:
-III-D Parameter Update
-The neural network parameters are learned by solving the following, unconstrained optimization problem:
+3.4 Parameter Update
+Neural network parameters are learned by solving an unconstrained optimization problem:
 where it should be understood that is short hand for all parameters of the neural network.
-The cost function is given by the following expression, where is a hyper-parameter to be tuned which controls regularization:
+The cost function is given by the following expression, where is a hyperparameter to be tuned which controls regularization:
 Parameters are updated iteratively using gradient-descent:
-where the hyper-parameter is the learning rate.
-In practice, some improved version of gradient-descent would be used instead, such as ADAM [24], but the former is easier to explain without loss of generality.
-The loss function is taken to be a modified version of the Least rquares Estimator (LSE), augmented with an extra term to account for partial derivatives:
-Note that two more hyper-parameters have been introduced: and .
-They act as weights that can be used to prioritize subsets of the training data in situations that call for it.
-By default, , meaning that all data is prioritized equally, which is the common use-case.
-However, in some situations, there may be an incentive to prioritize certain examples over others, as will be shown in the results section.
-This formulation offers the flexibility to do so.
-III-E Backward Propagation
-Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.s.t.) the inputs .
-By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
-III-E1 Loss Function Derivatives w.s.t. Neural Net Parameters
-In a first step, assume the quantities and are known and consider the gradient of the loss function with respect to the parameters from layer . Applying the chain rule to Eq. III-D yields:
-where the superscript has been dropped for notational convenience, with the understanding that all equations to follow apply to example . The various partial derivatives in Eq. 11 follow from Eq. 3:
-rubstituting these expressions into Eq. 11 yields:
-III-E2 Loss Function Derivatives w.s.t. to Neural Net Activations
+where the hyperparameter is the learning rate.
+In practice, some improved version of gradient-descent would be used instead, such as ADAM (kingma2017adam), but the former is easier to explain without loss of generality.
+The loss function is taken to be a modified version of the Least Squares Estimator (LSE), augmented with an extra term to account for partial derivatives:
+Note that two more hyperparameters have been introduced: and .
+They act as weights that can be used to prioritize subsets of the training data in situations that call for it (example later).
+By default, , meaning that all data is prioritized equally.
+3.5 Backward Propagation
+Forward propagation seeks to predict the gradient of outputs with respect to (abbreviated w.r.t.) inputs .
+By contrast, back-propagation computes the gradient of the cost function w.r.t.
+neural net parameters :
+3.5.1 Loss Function Derivatives w.r.t. Neural Net Parameters
+In a first step, assume the quantities and are known and consider the gradient of the loss function w.r.t.the parameters from layer . Applying the chain rule to Eq. 9 yields:
+where the superscript has been dropped for notational convenience, with the understanding that all equations to follow now apply to example .
+The various partial derivatives in Eq.
+11 follow from Eq.
+3:
+Substituting these expressions into Eq. 11 yields:
+3.5.2 Loss Function Derivatives w.r.t. to Neural Net Activations
 All quantities in Eq.
-15 are known, except for and which must be obtained recursively by propagating information backwards from one layer to the next, starting with the output layer .
+15 are known, except for and which must be obtained recursively by propagating information backwards from one layer to the next, starting with output layer .
 From Eq.
-III-D, it follows that:
+9, it follows that:
 The derivatives for the next layer are obtained through the chain rule.
 According to Eq.
-III-D, the loss function depends on activations and associated partials , which are themselves functions of and .
-Hence, according to the chain rule, current layer derivatives w.s.t.
-to previous layer activations are:
-Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds that:
-All necessary equations to compute the derivatives of the loss function w.s.t.
+9, the loss function depends on activations and associated partials , which are themselves functions of and .
+Hence, according to the chain rule, previous layer derivatives w.r.t.
+to current layer activations are:
+Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds:
+All necessary equations to compute the derivatives of the loss function w.r.t.
 parameters in any layer are now known.
@@
 17 after passing the solution in the current layer to the preceding layer as follows:
-III-E3 Cost Function Derivatives w.s.t. Parameters
-IV Verification and Validation
-The theory developed in the previous section was implemented into an installable Python library for which a link was provided earlier.
-This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig.
-2, where training data is shown as dots.
-These validation results compare JENN against a standard Neural Net (NN), defined as the equivalent neural network without gradient-enhancement.
-All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data, so long as the computational cost of obtaining the Jacobian is not prohibitive in the first place.
+3.5.3 Cost Function Derivatives w.r.t. Parameters
+Putting it all together, the derivative of the cost function w.r.t. neural network parameters are obtained by substituting Eq. 15 into Eq. 10, where the subscripts and are referenced to layer :
+A summary of key equations is available in A in vectorized form for computer programming, which have already been implemented in an open-source library called JENN developed for this work.
+4 Implementation
+An open-source Python library that implements JENN can be found online333https://pypi.org/project/jenn/.
+There exist many excellent deep learning frameworks, such as TensorFlow (tensorflow2015) and PyTorch (PyTorch), as well as many all-purpose machine learning libraries such as scikit-learn (sklearn2011).
+However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, which requires non-trivial effort to be implemented in those tools.
+The present library seeks to close that gap.
+It is intended for non-practitioners, who have access to partials and are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry to take advantage of them.
+JENN focuses exclusively on neural networks and serves as an upstream dependency to the popular Surrogate Modeling Toolbox (SMT) (saves2024smt) specializing in gradient-enhanced methods in general.
+4.1 Verification and Validation
+Using canonical test functions, Fig.
+2 compares JENN against a standard Neural Net (NN), taken to be the same neural network with gradient-enhancement turned off (i.e.
+).
+All else being equal, it can be seen that JENN outperforms NN in all cases, reaffirming the general concensus that gradient-enhancement achieves better accuracy with less data.
 In Figs.
@@
 In Fig.
-2d, where there are 100 samples, NN correctly captures the trend but contour lines are distorted, implying more samples would be needed to improve prediction.
-By contrast, under the same conditions, JENN predicts the Rastrigin function [25] almost perfectly.
-Finally, it was also verified that runtime scales as , as shown in Fig.
-2c.
-This plot was obtained using the Rastrigin function to generate progressively larger datasets, train a new model, and record runtime on a MacBook Pro (2021).
-V Example Application: Surrogate-Based Optimization
-Surrogate-based optimization is the process of using a fast-running approximation, such as a neural networks, to optimize some slow-running function of interest, rather than directly using the function itself.
-This use-case typically arises when the response of interest takes too long to evaluate relative to the time scale of the application.
-For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory[26].
-The purpose of this section is to quantify the benefit of JENN relative to regular neural networks for gradient-based optimization.
-This will be accomplished by solving the following optimization problem:
+2c, where there are 100 samples, NN correctly captures the trend but contour lines are distorted, implying more samples would be needed to improve prediction.
+By contrast, under the same conditions, JENN predicts the Rastrigin function (rastrigin1974systems) almost perfectly.
+The code to reproduce all results in this section is available online444code to reproduce available at https://codeocean.com/capsule/9360536/tree.
+4.2 Run Time
+It was empirically verified that runtime scales linearly as with sample size, as shown in Fig.
+3.
+This plot was obtained using the Rastrigin function (rastrigin1974systems) to generate progressively larger datasets, train a new model, and record runtime on a MacBook Pro (2021).
+4.3 Surrogate-Based Optimization
+Surrogate-based optimization is the process of using a fast-running approximation, such as a neural network, to optimize some slow-running function of interest, rather than directly using the function itself.
+This use case typically arises when the response of interest takes too long to evaluate relative to the time available for analysis.
+For example, in real-time optimal control, one may wish to consider a neural network to make quick predictions of complicated physics when updating the trajectory (Sanchez2018).
+The purpose of this section is to quantify the benefit of JENN relative to regular neural networks in the context of gradient-based optimization.
+This will be accomplished by solving the following, constrained optimization problem:
 | Minimize: | (21) | |||
@@
 | Subject To: | (23) | |||
-where is a surrogate-model approximation to the Rosenbrock [27] function, which is used as an optimization benchmark because it exhibits a shallow valley with near-zero slope, which creates a significant challenge for surrogate-based optimization:
-Results444code to reproduce available at https://codeocean.com/capsule/9360536/tree are shown in Fig.3b, where optimization histories are shown for the true response and three surrogate-model approximations: NN, JENN, and JENN (polished).
+where is a surrogate-model approximation to the Rosenbrock function (Rosenbrock1960), which is often used as a benchmark in optimization because the optimum lies in a “shallow valley” with near-zero slope that stalls convergence:
+Results are shown in Fig.4, which compares optimization convergence histories for the true response and three surrogate-model approximations: NN, JENN, and JENN (polished).
 Training data is not shown to avoid clutter, but was generated by sampling the true response on a regular grid plus 100 Latin hypercube points.
-The neural network training histories are also included in Fig.
-3a, in order to verify that both neural networks converged to the same degree.
-V-A Results Without Polishing
-Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not observed: neither JENN nor NN reach the true location of this optimum situated at .
-Carefully studying the contour lines of NN in the region where the Rosenbrock function flattens out, one can observe subtle undulations, within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
+4.3.1 Results Without Polishing
+Under perfect conditions, one would expect the optimization history to follow the same path for all models, but this is not the case.
+Neither JENN nor NN reach the true location of this optimum situated at with a value of zero.
+Carefully studying the contour lines of NN in the region where the Rosenbrock function flattens out, one can observe subtle undulations, likely within the noise of surrogate prediction error, which create artificial local minima that trap the optimizer.
+This behavior is the Achille’s heel of surrogate-based optimization methods.
 These effects are less pronounced for JENN, but the optimizer gets trapped nonetheless.
-In hindsight, this finding is not surprising because neural networks learn by minimizing prediction error (recall Eq.
-III-D).
-Hence, by construction, there is more reward during training for accurately predicting regions of large change, where prediction error implies a large penalty, than there is for accurately predicting flat regions in the vicinity of a zero optimum, where prediction error penalties are negligible.
+This finding is not surprising whenever the optimum lies in a flat region and approaches zero, since magnitudes are so small they cannot be distinguished from prediction error.
+For a broad set of engineering problems, the results attained by these baseline models may be sufficiently accurate.
+However, there are situations where optimization accuracy is desirable if not critical, such as aerodynamic shape optimization.
+In general, surrogate-based optimization tends to struggle any time the magnitude of the slopes in the region of interest lie within prediction noise.
 This situation is at odds with the needs of gradient-based optimization which requires extreme accuracy near the optimum; accuracy elsewhere is desirable but less critical, provided slopes correctly point in the direction of improvement.
-For a broad set of engineering problems, the results attained by these baseline models may be sufficiently accurate.
-However, there are situations where optimization accuracy is desirable if not critical, such as airfoil design where it is imperative that the final airfoil is free of undulations and non-smooth contours.
 An imperative research question therefore arises: how can these competing needs be reconciled?
-V-B Results With Polishing
-This situation can be ameliorated through “polishing.” Given a trained model, this approach consists of train it further after magnifying regions of interest.
+4.3.2 Results With Polishing
+This situation can be ameliorated through “polishing” which is uniquely afforded by this framework.
+Given a trained model, this approach consists of training it further after magnifying regions of interest.
 This is accomplished by judiciously setting to prioritize specific training points belonging to that subspace.
-For example, flat regions could be magnified by allocating more importance to small slopes during training, using a radial basis functions centered on such as:
+Specifically, flat regions can be magnified by allocating more importance to small slopes during training, using a radial basis functions centered on for example:
 where and are hyperparameters that control the radius of influence of a point.
@@
 Far away from any point where , the weights are whereas near those points .
-This has the desired effect of disproportionately emphasizing regions of small slope during training by up to a factor of 1000 at the expense of de-emphasizing other regions, which is why it is necessary to start from an already trained model.
-Final results are shown in Fig.
-3b for the Rosenbrock test case, where it can be seen that JENN (polished) now recovers the true optimum almost exactly.
-Overall, this section demonstrates that jacobian-enhancement offer a quantifiable benefit over standard neural nets for SBO applications.
-VI Noisy Partials
+This has the desired effect of disproportionately emphasizing regions of small slope during training up to a factor of 1000 at the expense of deprioritizing other regions, which is why it is necessary to start from an already trained model.
+Results are shown in Fig.
+4 for the Rosenbrock test case, where it can be seen that JENN (polished) now recovers the true optimum almost exactly.
+4.4 Noisy Partials
 We now consider the effect of noisy partials.
-As mentioned in the introduction, many practical design applications rely on legacy software suites, such as missile DATCOM [28] during early design phases, for which analytical or automatically differentiated partials are not available.
+Many practical design applications rely on legacy software suites during early design phases, such as missile DATCOM (DATCOM), for which analytical or automatically differentiated partials are not available.
 In this case, the only recourse for computing derivatives is finite difference.
-However, despite their attractive simplicity, it has been shown that finite difference methods are not as accurate as their counterparts [29].
-It is therefore natural to wonder: how accurate must partials be in order to take advantage of JENN in design applications where finite difference is unavoidable?
+However, despite their attractive simplicity, it has been shown that finite difference methods are not as accurate as their counterparts (MartinsHwang2013).
+It is therefore natural to wonder: how accurate must partials be in order to take advantage of JENN in design applications where finite difference might be unavoidable?
 The answer to this question is shown by example in Fig.
-4 using the Rastrigin function as a testbed555code to reproduce available at https://codeocean.com/capsule/9360536/tree once again, where corresponds to baseline results for a standard neural network (gradient-enhancement off).
-Training data partial derivatives were obtained using finite difference with progressively increasing steps sizes.
+5 using the Rastrigin function as a testbed once again, where corresponds to baseline results for a standard neural network (gradient-enhancement turned off).
+Partial derivatives of the training data were obtained using finite difference with progressively increasing step sizes.
 The corresponding mean finite difference error shown was computed against true analytical partials.
@@
 These results are welcome, as they demonstrate that the usefulness of gradient-enhanced methods is not restricted to perfectly accurate partials; even noisy partials can be useful (within a reasonable threshold).
-Finally, these results further show that there is no benefit to setting to values other than zero or one; it does not mitigate the effect of noisy partials.
-For most applications, the hyperparameter should therefore be treated as a binary scalar to turn gradient-enhancement on () or off (); except in special situations which call for distribution (rather than a scalar) such as polishing, as previously illustrated, or incomplete partials.
-The latter is likely to be a common scenario where partials are not available for all inputs, in which case can be used to turn gradient-enhancement on/off depending on whether a partial is missing or not (i.e.
-resulting in a tensor of zeros and ones).
-VII Conclusion
-In summary, this research developed a jacobian-enhanced, deep-learning formulation for dense multi-layer perceptrons and successfully validated the implementation against selected canonical test functions.
-It was shown that validation results support the general consensus that gradient-enhanced methods require less training data and achieve better accuracy than their non-gradient-enhanced counterparts, provided the cost of obtaining training data partials is not prohibitive in the first place.
-It was further shown that JENN outperforms standard neural networks in the context of surrogate-based optimization.
-However, it was observed that a learning procedure which minimizes any version of the least squares estimator is at odds with the needs of gradient descent in the vicinity of the optimum, where slopes are close to zero.
-It was further shown that this shortcoming can be overcome through polishing, which can be used to prioritize salient regions of interest through judicious use of hyperparameters.
-This unique feature enabled by the present formulation provides an additional knob for tuning optimized solution accuracy in situations that call for it.
-The ideas and findings of this paper have been published as an open-source library to allow readers to validate the results, and hopefully assist them in their surrogate modeling efforts.
-Acknowledgement
-The supporting code for this work used the exercises by Professor Andrew Ng [23] as a starting point.
-It then built upon it to include additional features but, most of all, it fundamentally changed the formulation to include jacobian-enhancement and made sure all arrays were updated in place (data is never copied).
-The author would therefore like to thank Professor Andrew Ng for offering the fundamentals of deep learning on Coursera, which masterfully explained what can be considered a challenging and complicated subject in an easy to understand form.
-VIII Key Equations in Vectorized Form
+These results further show that there is no benefit to setting to values other than zero or one; it does not mitigate the effect of noisy partials.
+For most applications, the hyperparameter should therefore be treated as a binary scalar variable to turn gradient-enhancement on () or off (); except in special situations such as incomplete partials or polishing, as previously illustrated, which call for a distribution of in that case.
+Incomplete partials refers to the scenario where partials are only available for some inputs, in which case can be used to turn gradient-enhancement off for the missing partials, i.e.
+resulting in a tensor of zeros and ones as illustrated next in airfoil application section.
+5 Application: Airfoil Shape Optimization
+This section quantifies the benefit of JENN for aerodynamic shape optimization through applied example.
+This is accomplished by optimizing the shape of a NACA 0012 airfoil in inviscid, transonic flow and comparing results to status quo neural networks.
+5.0.1 Problem Description
+To ensure reproducibility, the open-source software SU2 (Economon2016) was selected for computational fluid dynamics.
+The design problem is taken directly from their website555https://su2code.github.io/tutorials/Inviscid_2D_Unconstrained_NACA0012/, modified to hold lift coefficient fixed (instead of angle of attack), and airfoil geometry is parameterized using fourteen Hicks-Henne “bump” functions (Hicks1978).
+The goal is to minimize drag at some desired flight condition, specified by Mach number and lift coefficient :
+| Minimize: | (27) | |||
+| With Respect To: | (28) | |||
+| Subject To: | (29) |
+where denote the Hicks-Henne shape parameters.
+The flight condition parameters determine the function , which effectively defines a different instance of the optimization problem.
+Each instance took several minutes to solve using adjoint methods on a modern laptop.
+5.0.2 Statistical Regression
+Our goal is to accelerate numerical optimization by developing an optimal proxy, which can be used to near instantaneously shape an airfoil at any flight condition.
+In the proposed approach, this is accomplished by replacing the objective function by an equivalent approximation , taken to be a neural network.
+To do so, one must therefore sample a 16-dimensional input space composed of 14 shape variables and two flight conditions and to generate training data.
+Sample results are shown in Fig.
+6 for a subset of the design space.
+A total of 10,000 Latin hypercube samples were evaluated.
+Upon collecting samples, the data was used to determine how many training points were needed to achieve good prediction accuracy using JENN and regular Neural Networks (NN).
+This was accomplished by holding back a subset of the data for training and using the rest for testing.
+The neural net architecture (5 hidden layers of 16 neurons each) and hyperparameter settings were identical for both neural networks, except for gradient enhancement which was turned on or off accordingly.
+The results are shown in Fig.
+7 using test data to plot the R-squared value as function of the number of training samples.
+Results unambiguously show that JENN achieves the same prediction accuracy of 99% using five times fewer samples, re-affirming the conclusion by other authors that gradient-enhanced methods reduce the amount of data required for good prediction, even on practical engineering problems.
+Using the results of Fig.
+7, it is possible to identify the minimum number of points needed to train an accurate neural net under both paradigms for this problem, as indicated by the arrows.
+Associated convergence properties and goodness of fit were verified to be satisfactory but not reported for conciseness, except for Table 1 which compares the R-squared value of the partials in order to emphasize that JENN predicts them more accurately.
+| NN | -0.69 | 0.79 | 0.79 | 0.68 | 0.32 | 0.26 | 0.32 | -0.95 | 0.73 | 0.80 | 0.77 | 0.59 | 0.49 | -0.10 |
+| JENN | 0.37 | 0.90 | 0.89 | 0.78 | 0.59 | 0.18 | 0.34 | 0.29 | 0.85 | 0.88 | 0.87 | 0.77 | 0.59 | 0.13 |
+Corresponding sensitivity profiles are shown in Fig.
+8 for all variables, in order to show the topological differences that resulted from gradient enhancement.
+Overall, it was found that JENN and NN exhibit excellent agreement in the subsonic regime, but that discrepancies start to appear in transonic flow at higher Mach numbers, where shock waves are present, presumably creating stronger sensitivity to shape variables that control the region where shocks tend to appear.
+As a result, the surrogate model learns to predict sensitivities more accurately in some regions than others, which can be seen from the prediction profiles.
+Although there is good agreement in general, notice that NN slopes occasionally point in a different direction in regions that tend to be flatter.
+This could have subtle consequences during optimization.
+Finally, it should be noted that SU2 did not provide partial derivatives for or , which implies incomplete gradient information.
+This is particularly noteworthy because most engineering design applications will fall in this scenario: partials are only available for some variables.
+Yet, by virtue of the framework afford by JENN, such missing information is easily overcome by simply setting for the missing partials.
+This represents a simple but important practical improvement over similar reported work.
+5.0.3 Surrogate-Based Optimization
+The surrogate models developed in the previous subsection were substituted for in the optimization problem described by Eqs.
+27–29, which was solved using the SLSQP algorithm from Scipy (2020SciPy-NMeth).
+Fig.
+4 compares final solutions against the initial geometry using different approaches.
+All solutions successfully recovered a supercritical airfoil shape, as expected, but adjoint-based optimization took several minutes, whereas surrogate-based optimization took less than a second, setting the cost of generating training data aside.
+However, there are some observable differences.
+In particular, notice the presence of a weak but unwanted secondary shock wave in the answer obtained using a regular NN.
+This is necessarily attributed to small differences in predicted partial derivatives, since that is the only difference between JENN and a regular NN.
+This underlines the foreseeable conclusion that gradient-enhanced methods ameliorate surrogate-based optimization somewhat whenever gradient-descent algorithms are used.
+Overall, predictions remained comparable, but the cost of generating data was significantly less for JENN.
+6 Conclusion
+In summary, this research developed JENN, which are densely connected multi-layer perceptrons that predict gradients more accurately, successfully validated the implementation against canonical test functions, and applied it to a practical airfoil shape optimization problem.
+All results support the general consensus that gradient-enhanced methods require less training data than their status quo neural network counterparts and achieve better accuracy, provided the cost of obtaining training data partials is not prohibitive in the first place.
+It was further shown that JENN outperforms standard neural networks in the context of surrogate-based optimization, but that surrogate-based methods are more prone to getting trapped in artificial local minima, whenever the true optimum lies in flat regions and approaches zero.
+This was apparent for the Rosenbrock test function, but also for the airfoil design problem where the regular NN solution exhibited a weak secondary shockwave that was not present with JENN.
+Overall, JENN tends to achieve optimal solutions that are in better agreement with the true optimum.
+Furthermore, through a process dubbed polishing, uniquely afforded by this framework, JENN enables salient regions of interest to be magnified through judicious use of hyperparameters.
+This feature provides an additional knob for improving surrogate-based optimization accuracy in situations that call for it, as demonstrated with the Rosenbrock example.
+The same mechanism also enables missing partials to be ignored during training by setting for those partials, such that gradient-enhancement can be leveraged even in situations when information is incomplete.
+This is likely to be a common scenario for many engineering design problems, as was the case for the airfoil design example where and were missing.
+This simple mechanism is likely to be of important practical value.
+In closing, despite its advantages over status quo neural networks, JENN does not overcome the curse of dimensionality; although it alleviates it by reducing the number of samples needed.
+Sampling remains necessary to generate training data, and so the practical utility of gradient-enhanced methods is likely to be restricted to problems where intrinsic dimensionality is either small to start with or can be recovered using elaborate dimensionality reduction techniques (BerguinMavris2015; BerguinRancourtMavris2015; Constantine2015).
+To be adopted at industrial scales, new techniques are needed to avoid high-dimensional sampling whilst retaining the benefit of optimal proxies for accelerating numerical optimization.
+Acknowledgments
+The supporting code for this work used the exercises by Professor Andrew (Ng2017) as a starting point, but fundamentally change the formulation to handle Jacobian enhancement.
+The primary author, who was then a student, would therefore like to thank Professor Ng for offering the fundamentals of deep learning on Coursera, which masterfully explained what can be considered a challenging and complicated subject in an easy way to understand.
+Appendix A Key Equations in Vectorized Form
 This section provides vectorized equations to facilitate re-implementation or as a companion guide to the JENN library.
-VIII-A Data Structures
+A.1 Data Structures
 Using capital letters, notation can be extended to matrix notation to account for all training examples, starting with training data inputs and outputs:
@@
 Finally, the back propagation derivatives w.r.t. parameters are:
-VIII-B Forward Propagation (Eq. 3)
-VIII-C Parameter Update (Eq. 8)
-VIII-D Backward Propagation (Eqs. 15–16, 17, 20)
+A.2 Forward Propagation (Eq. 3)
+A.3 Parameter Update (Eq. 8)
+A.4 Backward Propagation (Eqs. 15–16, 17, 20)
 The loss function partials w.r.t. last layer activations are: